OpenStudy (anonymous):

Input the equation of the given line in standard form. The line including the two points (5, 3), (8, -2)

OpenStudy (anonymous):

Standard form is Ax + By + C = 0 first put it into y = mx + b form, which in this case is y = -(5/3)x + 34/3 move your x over to the left (5/3)x + y = 34/3 multiply through 3 to clear the fractions 5x + 3y = 34 5x + 3y - 34 = 0